Reports 1-1 of 1 Clear search Modify search
IOO (Laser Bench) - 23:27 Wednesday 19 December 2018 (7451) Print this report
Installation of High power beam shutter

Kokeyama-san, Ge Guiguo

Today we remove the old beam shutter and install the high beam shutter(without water cooler beam dampFIG1, without the cover because lack of proper M2 screws), Old beam shutter is return to shutter box near MCE chamber(FIG4).
The new shutter driver  need +-24V/0.6A power(FIG.3) and we set shutter driveri under the optical table near the laser.(FIG2).
Now with the new beam shutter we can locked IMC.

But now the high power beam shutter is open at 0V that is unsafety(old shutter is close at 0V),and it also mean the setting of MEDM and guardian is oppsite to the previous, so we modify the medm of PMC and IMC temporarily.(FIG.5.6 is medm of PMC), I will check the guardian tomorrow.
We will change the solenoid  wihich can close the shutter at 0V in the near future, after that we need modify the medm and guardian again.

Images attached to this report
Comments to this report:
keiko.kokeyama - 10:25 Thursday 20 December 2018 (7456) Print this report
imc guardian was already modified for the current shutter.
Search Help